### v1

# How to set AdamW’s weight decay as you scale model and dataset size

Abstract
We show that weights learned by AdamW can be understood as an exponential moving average (EMA) of recent updates. This gives critical insights for how to set the weight decay in AdamW, and how the weight decay should scale with model and dataset size. In particular, the key hyperparameter for an exponential moving average is the EMA timescale. Intuitively, the EMA timescale can be understood as the number of recent iterations the EMA averages over. Given a fixed learning rate, there is a one-to-one mapping from the EMA timescale to the usual weight decay hyperparameter. Thus, choosing an EMA timescale implicitly sets the weight decay. Importantly, there are natural guidelines for sensible values for the EMA timescale: we need to average over all datapoints, so the EMA timescale should not be (much) smaller than 1 epoch, and we need to forget early updates, so the EMA timescale should not be (much) bigger than the total number of training epochs. In our experiments, we find that optimal EMA timescales are consistent with these guidelines, as are the hyperparameters chosen in recent large-scale LLM pretraining runs (e.g. Llama 1+2 and Stable LM). Critically, these guidelines suggest that the optimal EMA timescale should not change (much) as we scale the model and dataset. That implies that as the dataset size increases, the optimal weight decay should fall. Moreover, as the model size increases, the optimal weight decay should also increase (if we follow the muP recommendation for scaling the learning rate).
1 Introduction
A common machine learning workflow is to prototype by training many smaller models, then at the end to do one final training run, with the largest possible model on the largest possible dataset. This workflow is used in many settings, from small scale student projects to the largest LLM training runs. However, for this workflow to be effective, we need to understand how to transfer optimal hyperparameters from smaller-scale prototyping runs to the final, largest model and dataset. In this paper, we focus on transferring the weight decay hyperparameter in AdamW (Loshchilov & Hutter, 2018), as AdamW is used in many settings including the largest LLM pretraining runs (e.g. Zhang et al., 2022; Touvron et al., 2023a, b; Tow et al., 2023).
The most obvious approach for changing the hyperparameters as we scale the model size (but not the dataset size) is µP (Yang et al., 2022). The key recommendation from µP is to scale the learning rate, $\eta$, as $1/\text{fan\_in}$. However, the theory behind µP analyses only how the learning rate affects behaviour close to initialization as we scale model size; this theory says nothing about the effect of regularization as the optimizer approaches convergence. As such, µP does not provide the whole answer to hyperparameter transfer in AdamW, given the need to additionally transfer weight decay hyperparameters, $\lambda$, across model and dataset sizes.
To understand how to set the AdamW weight decay and how to transfer it across model and dataset sizes, we show that AdamW can be understood as an exponential moving average (EMA). Of course, both Adam and AdamW use EMAs to estimate the average gradient, $m_{t}$, and average squared gradient, $v_{t}$. That is not what we are talking about. Instead, we observe that in algorithms with decoupled weight decay (i.e. AdamW but not Adam), the weights themselves are an EMA of recent updates (see Sec. 3.1 for details).
The key hyperparameter in an EMA is the EMA timescale, which intuitively describes the number of previous iterations that the EMA averages over; this timescale is $\tau_{\text{iter}}=1/(\eta\lambda)$ (Sec. 3.1). Thus, setting the learning rate, $\eta$, and timescale, $\tau_{\text{iter}}$, implies a value for the weight decay, $\lambda=1/(\eta\tau_{\text{iter}})$. As such, given a fixed learning rate, $\eta$, we can implicitly set the weight decay, $\lambda$, by setting the EMA timescale, $\tau_{\text{iter}}$. We argue that it is easier to reason about and tune the EMA timescale, $\tau_{\text{iter}}$, because there is a natural range of values for $\tau_{\text{iter}}$. Specifically, if we measure the timescale in epochs, i.e. $\tau_{\text{epoch}}=\tau_{\text{iter}}/M$, where $M$ denotes the number of iterations per epoch, then $\tau_{\text{epoch}}$ should not be far smaller than 1, otherwise the EMA would not average over all the datapoints. At the same time, the $\tau_{\text{epoch}}$ should not be far larger than the total number of training epochs, otherwise the algorithm would be unable to forget contributions to the EMA from very early on in the training process.
We perform a number of experiments showing this natural range for the optimal $\tau_{\text{epoch}}$ seems to capture the actual optimal $\tau_{\text{epoch}}$ (Sec. 4.1). Moreover, we note that this natural range of values has been used for successful large-scale LLM training runs such as Llama 1+2 (Touvron et al., 2023a, b) and Stable LM 3B (Tow et al., 2023).
Next, we notice that this natural range of values for $\tau_{\text{epoch}}$ does not seem to change with model and dataset size. That has important implications for how to scale the weight decay with the model and dataset size. Specifically, as the dataset size increases, we find that the optimal weight decay falls (Sec. 4.2). Moreover, if we follow the usual µP recommendation for the learning rate as we increase the model sizes, we find that the optimal weight decay should increase as the model size increases (Sec. 4.3). We validate both of these predictions experimentally.
2 Background
Yang et al. (2022) considered how to transfer the learning rates in SGD and Adam across model sizes. They considered two key desiderata. First, the initial random weights, $W^{0}\in\mathbb{R}^{D\times\text{fan\_in}}$, times an input, $x\in\mathbb{R}^{\text{fan\_in}}$, should not grow or shrink as the model changes size. We could write this requirement as, $W^{0}x\sim 1$, where we use $\sim$ to indicate “scales with”. This gives the usual $W^{0}\sim 1/\sqrt{\text{fan\_in}}$ scaling of the standard deviation of the initial random weights. At the same time, they required that the change in the outputs, caused by the first weight update, $\Delta Wx$, shouldn’t grow or shrink as the model changes size; i.e. $\Delta Wx\sim 1$. Yang et al. (2022) show that this requirement implies that the learning rate should scale as $1/\text{fan\_in}$.
The distinction between the $1/\sqrt{\text{fan\_in}}$ scaling of the initial weights vs. the $1/\text{fan\_in}$ scaling of the learning rate might be a bit puzzling. The intuitive reason for this distinction is given in Yang et al. (2022) Appendix J. In short, for the initial weights,
$y_{i}=\sum_{j=1}^{\text{fan\_in}}W^{0}_{ij}x_{j}$ (1)
we know that $y_{i}\sim\sqrt{\text{fan\_in}}\times W^{0}_{ij}$, so to ensure that $y_{i}\sim 1$, we need $W^{0}_{ij}\sim 1/\sqrt{\text{fan\_in}}$. However, this square root scaling only arises in very specific circumstances, when each term in the sum, $W^{0}_{ij}x_{j}$, is zero-mean and uncorrelated. These requirements hold for the initial weights, as they are sampled IID from a zero-mean distribution. However, these conditions do not hold for the update (the precise reason why is complex; see Yang et al., 2022, for details),
$\Delta y_{i}=\sum_{j=1}^{\text{fan\_in}}\Delta W_{ij}x_{j},$ (2)
thus, $\Delta y_{i}\sim\text{fan\_in}\times\Delta W_{ij}$. To ensure that $\Delta y_{i}\sim 1$, we therefore need $\Delta W_{ij}\sim 1/\text{fan\_in}$. As the Adam updates, $\Delta W_{ij}$ scale with the learning rate, $\eta$, this implies $\eta$ must also scale with $1/\text{fan\_in}$.
To see why Adam updates scale with the learning rate, consider an Adam update,
$\Delta W_{ij}$ $=\eta\frac{\hat{m}_{ij}}{\sqrt{\hat{v}_{ij}}}$ (3)
where we have neglected the small $\epsilon$. Here, $\hat{m}_{ij}$ is an EMA estimate of the gradient $g_{ij}$, while $\hat{v}_{ij}$ is an EMA estimate of the expected squared gradient $g_{ij}^{2}$. Thus, $\hat{m}_{ij}\sim g_{ij}$ and $\hat{v}_{ij}\sim g_{ij}^{2}$. That implies $\hat{m}_{ij}/\sqrt{\hat{v}_{ij}}\sim 1$, so looking back at Eq. 3, we have $\Delta W_{ij}\sim\eta$. Hence, to get $\Delta W_{ij}\sim 1/\text{fan\_in}$, we need $\eta\sim 1/\text{fan\_in}$, which is the origin of the $1/\text{fan\_in}$ scaling of the learning rates in µP.
Importantly, as Yang et al. (2022) considers the size of the first update relative to the random initial weights, it does not give guidance about how to change the weight decay with the model size, as the weight decay only becomes relevant after many learning steps.
3 Methods
3.1 AdamW as an EMA
An AdamW update for a single parameter at the $t$th iteration can be written as,
$w_{t}$ $=(1-\eta\lambda)w_{t-1}-\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (4)
where $w_{t}$ is a neural network weight, $\eta$ is the learning rate, $\lambda$ is the weight decay, $\epsilon$ is a small constant, $\hat{m}_{t}$ is a bias-correct EMA estimate of the expected gradient, and $\hat{v}_{t}$ is a bias-correct EMA estimate of the expected squared gradient. Notice that here we adopt the parameterization used by major optimization libraries such as torch.optim [^1] https://pytorch.org/docs/stable/generated/torch.optim.AdamW.html and optax [^2] https://optax.readthedocs.io/en/latest/api/optimizers.html#optax.adamw, where the decay for $w_{t-1}$ is controlled by the product $\eta\lambda$. This parameterization is different from the form suggested in the original AdamW paper (Loshchilov & Hutter, 2018), which we will discuss in detail in the related work section.
Now we will show that these weight updates (Eq. 4) can be understood as an EMA. Recall that a generic exponential moving average estimate, $\text{ema}_{t}$ can be written as,
$\operatorname{ema}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})\operatorname{ema}_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\;q_{t}.$ (5)
Here, $\operatorname{ema}_{t}$ forms an exponential moving average estimate of $q_{t}$, and the EMA timescale is $\tau_{\text{iter}}$. Specifically, if we take the canonical form for an EMA (Eq. 5) and set:
$\nicefrac{{1}}{{\tau_{\text{iter}}}}$ $=\eta\lambda$ $\operatorname{ema}_{t}$ $=w_{t}$ $q_{t}$ $=-\frac{1}{\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (6)
then we recover the AdamW updates (Eq. 4). Of course, AdamW uses EMAs to compute $\hat{m}_{t}$ and $\hat{v}_{t}$. Our key insight was that additionally, the overall AdamW updates for $w_{t}$ can themselves be understood as EMAs.
Importantly, there are only two hyperparameters under the EMA view: the EMA timescale, $\tau_{\text{iter}}=1/(\eta\lambda)$ and the initialization, $\operatorname{ema}_{0}$. The connection between AdamW and EMA thus suggests that AdamW also should have only two hyperparameters, $\tau_{\text{iter}}$, and something about the initialization. In Appendix A we show just this. Specifically, for scale-invariant models, the full trajectory of an AdamW optimizer depends only on $\tau_{\text{iter}}$, and the relative size of the parameter initialization against the learning rates.
3.2 Implication of the EMA interpretation
The previous section argued that AdamW is an EMA, and the key hyperparameter in an EMA is the timescale, $\tau_{\text{iter}}$. Here, we argue that the EMA viewpoint suggests a range of sensible values for the EMA timescale. We can rewrite AdamW’s EMA as a weighted average of all updates (see Appendix B for a derivation of this form).
$w_{t}$ $=\nicefrac{{1}}{{\tau_{\text{iter}}}}\sum_{t^{\prime}=1}^{t}e^{-(t-t^{\prime})/\tau_{\text{iter}}}\left(-\frac{1}{\lambda}\frac{\hat{m}_{t^{\prime}}}{\sqrt{\hat{v}_{t^{\prime}}}+\epsilon}\right)$ (7)
where we make a particular choice of the (unnormalized) EMA weights, $e^{-(t-t^{\prime})/\tau_{\text{iter}}}$ . These weights highlight that recent updates (i.e. $t^{\prime}$ close to $t$) contribute more to the weighted average. Specifically, for recent updates where $t-t^{\prime}$ is far smaller than $\tau_{\text{iter}}$, the weights, $e^{-(t-t^{\prime})/\tau_{\text{iter}}}$, are around $1$. At the same time, updates from further into the past contribute less: when $t-t^{\prime}$ is far larger than $\tau_{\text{iter}}$, then the weights, $e^{-(t-t^{\prime})/\tau_{\text{iter}}}$, decay to zero. Critically, this implies that the EMA averages over roughly the last $\tau_{\text{iter}}$ minibatch updates.
However, it is still difficult to understand the implications of averaging over the previous $\tau_{\text{iter}}$ iterations. Instead, we propose to work with the timescale measured in terms of epochs,
$\tau_{\text{epoch}}=\tau_{\text{iter}}/M.$ (8)
where $M$ is the total number of minibatches in the dataset or equivalently iterations in an epoch.
Now, $\tau_{\text{epoch}}$ has a natural interpretation. Specifically, if $\tau_{\text{epoch}}$ is around $1$, then the AdamW EMA averages over updates in the last epoch. Alternatively, if $\tau_{\text{epoch}}$ is far smaller than $1$, then the AdamW EMA averages over only a small fraction of the training samples, which is likely to be very suboptimal. Likewise, if $\tau_{\text{epoch}}$ is too big (compared to the total number of epochs in the full training run), AdamW will be unable to forget the random initialization and/or very early gradients, which again could lead to suboptimal performance.
4 Results
In this section, we first demonstrate that the optimal value of $\tau_{\text{epoch}}$ falls within a natural range. We then show that as dataset sizes vary, the optimal $\tau_{\text{epoch}}$ remains within a narrow range for the same model. Finally, we illustrate that across different model sizes, the optimal timescale also remains consistent within a close range.
For all experiments, we implemented the model using PyTorch and performed optimization using its AdamW implementation. For all tasks, we did not use weight decay on the normalization layers. We used a cosine learning rate schedule, which reduced the learning rate by a factor of 10. All experiments were conducted on an internal cluster of Nvidia 1080ti/2080ti GPUs. Additional computation was expended on preliminary experiments. The results presented are the averages from three distinct random seeds. Comprehensive details on the hyperparameter search range and model specifications can be found in Appendix E.
4.1 The natural range of values for the AdamW EMA timescale
We considered a ResNet-18 (He et al., 2016) and ViT (Dosovitskiy et al., 2021) trained on CIFAR-10 (Krizhevsky, 2009) under a batchsize of 100, as such we have $M=50,000/100=500$. We trained the models with initial learning rates, $\eta$ of $\{10^{-4},10^{-3},10^{-2}\}$ for 200 epochs. For each $\eta$, we sweep $\tau_{\text{iter}}=1/(\eta\lambda)$ from $10^{-7}$ to $10^{-3}$ by changing $\lambda$.
The results are presented in Fig. 1: Broadly, we found that the optimal $\tau_{\text{epoch}}$ was usually between 1 and 200, which is clearly within our guidelines. While the optimal timescale was higher for the ViT with a learning rate of $10^{-4}$, the relevance of this finding is unclear as this learning rate is clearly suboptimal.
Figure 1: Optimal EMA timescale $\tau_{\text{epoch}}$ lies in a natural range. We trained a ResNet-18 (A) and a ViT (B) on CIFAR-10 using AdamW with cosine learning rate decay schedule under various initial learning rates $\eta$ (different lines) and different $\tau_{\text{epoch}}$ (x-axis; note we set $\lambda=1/(\eta M\tau_{\text{epoch}})$). We plotted various performance metrics after 200 epochs. The optimal values in terms of test metrics are highlighted with red crosses: For both models, the optimal $\tau_{\text{epoch}}$ lies in a natural range between 1 and 200 (shown by the gray dashed lines), except for $\eta=10^{-4}$ for ViT, which shows clearly suboptimal performance.
Moreover, we looked at large-scale LLM pretraining runs (Table 1) from Llama 1 (Touvron et al., 2023a), Llama 2 (Touvron et al., 2023b) and Stable-LM (Tow et al., 2023). Large-scale LLM pretraining runs are particularly interesting in our context because they are usually trained for only 1 or a few epochs. In that context, the EMA viewpoint suggests a narrow range of optimal weight decays: that $\tau_{\text{epoch}}$ should be neither much smaller nor much larger than $1$. While the use of a schedule, with smaller timescales initially and longer timescales at the end does complicate things slightly, the practical values chosen accord with those suggested by the EMA viewpoint. In particular, the initial timescales are smaller than 1 epoch, to ensure that very early contributions to the EMA may be forgotten, while the final timescale is around the total number of epochs, to ensure averaging over all datapoints. To obtain these timescales, we used the fact that that $\lambda=0.1$ for Llama 1, Llama 2, Stable LM, and Llama 1 + 2 decay the learning rate by a factor of 10, while Stable-LM decays to 4% of initial learning rate (Touvron et al., 2023a, b) and Stable-LM (Tow et al., 2023).
Table 1: EMA timescales in large-scale LLM pretraining runs
Model initial lr tokens/batch tokens epochs init $\tau_{\text{epoch}}$ final $\tau_{\text{epoch}}$
Llama 1 6.7B + 13B $3\times 10^{-4}$ 4 M 1 T 1 0.133 1.33
Llama 1 33B + 65B $1.5\times 10^{-4}$ 4 M 1.4 T 1 0.190 1.90
Llama 2 6.7B + 13B $3\times 10^{-4}$ 4 M 2 T 1 0.067 0.67
Llama 2 34B + 70B $1.5\times 10^{-4}$ 4 M 2 T 1 0.133 1.33
Stable-LM 3B $3.2\times 10^{-4}$ 4 M 1 T 4 0.125 3.13
4.2 Transferring weight decay across dataset sizes
Next, we study the relationship between optimal $\tau_{\text{epoch}}$ and the dataset size. Our prediction is that as dataset size increases, the optimal $\tau_{\text{epoch}}$ should remain broadly fixed. If we increase the dataset size with minibatch size fixed, then the number of minibatches in the dataset/an epoch, $M$, increases. Rearranging Eq. (8), that implies that $\tau_{\text{iter}}$ increases with dataset size, $\tau_{\text{iter}}=M\tau_{\text{epoch}}$. And rearranging $\tau_{\text{iter}}=1/(\eta\lambda)$,
$\lambda$ $=\frac{1}{\eta\tau_{\text{iter}}}=\frac{1}{\eta M\tau_{\text{epoch}}},$ (9)
tells us that if the optimal $\tau_{\text{epoch}}$ changes a little as the dataset size increases, we would predict that the optimal $\lambda$ should decrease as the dataset size increases.
To confirm this hypothesis, we trained a ResNet-18 and a ViT on ImageNet (Fig. 2). We used the $32\times 32$ downscaled version of ImageNet provided by Chrabaszcz et al. (2017) to ensure that we were able to perform the large number of training runs required to assess performance for different hyperparameter settings and dataset sizes. We trained the models on different subsets of ImageNet, where we randomly drew $80,160,320$, and $640$ samples from each of the $1,000$ classes except for the largest subset of $1.28M$ samples in total, where we randomly drew a subset from the whole dataset. We used a fixed batch size of 100 for all runs. We used an initial learning rate of $10^{-3}$ and swept $\lambda$ between $10^{-3}$ to $10^{3}$ and then plotted the performance metrics after 50 epochs. In the top row of Fig. 2A and Fig. 2B, we plotted the performance metrics vs. $\lambda$. Note that the optimal $\lambda$ decreases dramatically as the dataset size increases. In the bottom row of Fig. 2A and Fig. 2B, we plotted the performance metrics vs. $\tau_{\text{epoch}}$. Critically, the optimal $\tau_{\text{epoch}}$ is far more stable than the optimal $\lambda$.
Figure 2: The optimal $\tau_{\text{epoch}}$ transfers across dataset sizes. We trained ResNet-18 (A) and ViT (B) on subsets of downsampled ImageNet of various sizes (lines of different colors) under different weight decay (dots on the lines) under a fixed batch size of 100. An initial learning rate of $10^{-3}$ is used with cosine decay scheduling. The performance metrics after 50 epochs are plotted against the weight decay $\lambda$ and the corresponding timescale $\tau_{\text{epoch}}$ computed with the initial learning rate. The dashed lines show the optimal $\tau_{\text{epoch}}$ at a subset size of $320,000$: In both models, the optimal $\tau_{\text{epoch}}$ is fairly stable across dataset sizes whereas the optimal $\lambda$ decreases dramatically as dataset size grows.
4.3 Transferring weight decay across model sizes
Figure 3: The optimal $\lambda$ increases with model size whereas the optimal timescale is more stable. We trained ResNet-18 on a subset of ImageNet 32x32, with varying width factor $s$ (lines of different colors) under a fixed base learning rate $10^{-3}$ with varying weight decay (dots on the lines) and plotted the metrics after 50 epochs vs. weight decay strength. The top row scales the hyperparameters using the direct µP approach (Eq. 11; i.e. fixed $\lambda$), while the bottom row scales the hyperparameters to ensure $\tau_{\text{iter}}$ is fixed (Eq. 13; $\lambda$ increases with model size). Note that as $\eta_{\text{base}}=10^{-3}$ is fixed, there is a direct relationship between the optimal $\lambda_{\text{base}}$ and the optimal $\tau_{\text{iter;base}}$.
Figure 4: AdamW breaks the learning rate scaling of µP. Following the experiment setting in Yang et al. (2022), we trained a ResNet-18 with varying width factor $s$ (lines of different colors) under various base learning rates $\eta_{\text{base}}$ (x-axis) on CIFAR-10 (A) and a $320,000$ samples subset of ImageNet 32x32 (B). We then plotted the metrics after 200 (for CIFAR-10) and 50 (for ImageNet) epochs against $\eta_{\text{base}}$. The top row scales the hyperparameters using the direct µP approach (Eq. 11; i.e. fixed $\lambda$), while the bottom row scales the hyperparameters to ensure $\tau_{\text{iter}}$ is fixed (Eq. 13; $\lambda$ increases with model size). In both datasets, the direct approach breaks the stability of optimal $\eta_{\text{base}}$ in terms of test metrics due to changing the timescale whereas our scaling allows for consistent $\eta_{\text{base}}$ across model sizes.
Of course, it is also critical to understand how to modify the AdamW hyperparameters as we increase the model size. The most obvious approach to this issue is µP (Yang et al., 2022), which predicts that the optimal learning rate should decrease as $1/\text{fan\_in}$, by considering the behavior of the first few learning steps relative to the random initial weights. However, the theory behind µP did not consider how weight decay affects the learned weights in the later phase of the optimization.
Formally, we take $\eta_{\text{base}}$ and $\lambda_{\text{base}}$ as the learning rate and weight decay tuned on a smaller base model, with fan-in of $\text{fan\_in}_{\text{base}}$. Now, we consider scaling the model width by a factor of $s$,
$s$ $=\frac{\text{fan\_in}}{\text{fan\_in}_{\text{base}}}$ (10)
The most direct approach using µP scaling with AdamW (e.g. used by Lingle, 2024) scales the learning rate as recommended by µP, while keeping the weight decay fixed,
$\eta$ $=\eta_{\text{base}}/s$ $\lambda$ $=\lambda_{\text{base}}.$ (11)
However, if we look at the implied EMA timescale,
$\tau_{\text{iter}}$ $=1/(\eta\lambda)=s/(\eta_{\text{base}}\lambda_{\text{base}})=s\;\tau_{\text{iter;base}}$ (12)
we see that this approach scales the EMA timescale with the model size. However, one of our hypotheses is that the optimal timescale should not vary with the model size. As such we conjecture that the scaling in Eq. (12) would break the transferability of optimal learning rate across model sizes.
How do we resolve this issue? We have two desiderata. First, from µP, we should scale $\eta$ with $1/\text{fan\_in}$. Second, we have the timescale $\tau_{\text{iter}}=1/(\eta\lambda)$ fixed. It might at first seem difficult to reconcile these two desiderata. Indeed, it is impossible in the usual setting where $\lambda$ is fixed. However, it is possible to reconcile them if we allow the weight decay, $\lambda$, to strengthen for larger models.
$\eta$ $=\eta_{\text{base}}/s$ $\lambda$ $=s\;\lambda_{\text{base}}.$ (13)
Thus,
$\tau_{\text{iter}}$ $=1/(\eta\lambda)=1/(\eta_{\text{base}}\lambda_{\text{base}})=\tau_{\text{iter;base}}.$ (14)
So the EMA timescale remains fixed as the model size increases.
We begin by testing our prediction that the optimal EMA timescale was constant across model sizes, while the optimal weight decay was not, we trained a ResNet-18 of varying widths using µP’s codebase on $320,000$ samples subset of downscaled ImageNet, where we randomly drew $320$ samples from each class. We scaled the models’ widths by factors of 2, giving $s\in\{0.5,1,2\}$. We fixed $\eta_{\text{base}}=10^{-3}$ and used $\eta=\eta_{\text{base}}/s$ to change the learning rate with network size. We swept $\lambda_{\text{base}}$ from $10^{-2}$ to $10^{1}$. We then considered two ways of modifying the weight decay as we scaled the network and plotted the metrics after 50 epochs against weight decay. Fig. 3 (top row) leaves $\lambda$ fixed as the network size increases (Eq. 11). We can see the optimal $\lambda_{\text{base}}$ varies dramatically across network sizes. In contrast, Fig. 3 (bottom row) increases the weight decay as the network size increases (Eq. 13), as that leaves $\tau_{\text{iter}}$ unchanged (Eq. 14). It is evident that the optimal $\lambda_{\text{base}}$ and hence $\tau_{\text{iter}}$ is far more stable in this setting. Notably, while e.g. the optimal $\lambda_{\text{base}}$ for test loss on the bottom row does appear different from the others, if anything, this would indicate that the relationship between $s$ and $\lambda$ is even stronger than expected. In particular, this point indicates that even as we scale $\lambda$ with $s$, it the optimal $\lambda_{\text{base}}$ may still be increasing with $s$. That would seem to indicate that the relationship between $s$ and $\lambda_{\text{base}}$ might be superlinear (e.g. $\lambda_{\text{base}}\propto s^{\alpha}$, where $\alpha>1$), though we would need more, larger-scale experiments to definitively establish any such super-linearity, which is out-of-scope for the present work.
Importantly, scaling the weight decay correctly has important implications even for the original µP predictions about how the optimal learning rate transfers across model sizes. Specifically, we trained a ResNet-18 on CIFAR-10 and again, on the 320K downscaled ImageNet subset, with a fixed $\lambda_{\text{base}}=1.0$ and swept $\eta_{\text{base}}$ from $10^{-5}$ to $10^{-2}$. In the top row of Fig. 4A and B, we use the direct µP scaling (Eq. 11), with a fixed weight decay. Remarkably, we found that the optimal base learning rate varied dramatically across model sizes. This indicates that µP scaling of the optimal learning rate breaks down in AdamW. Indeed, the original µP paper (Yang et al., 2022) did not examine AdamW, and this finding is confirmed by recent work on hyperparameter transfer in large-scale transformers (Lingle, 2024).
We hypothesized that this breakdown was due to the timescale changing with model size (Eq. 12), and thus, that we could restore the usual µP scaling of the learning rate by increasing the weight decay with model size (Eq. 13). To confirm this finding, in the bottom row of Fig. 4A and B, we increase the weight decay with model size (Eq. 13). We found that the optimal base learning rate was now far more stable when varying model sizes, confirming our hypothesis. Similar behavior was also observed when training ViTs of different widths on the ImageNet subset, the results are presented in Appendix C.
Finally, the size of the weight decay has implications for the magnitude of the learned weights, and hence whether the outputs $Wx$ remain $\mathcal{O}\left(1\right)$ which was one of the key desideratum used to derive µP (Yang et al., 2022). In particular, in AdamW the magnitude of the learned parameters, $W_{ij}$ scales with $1/\lambda$ (or $W_{ij}\sim 1/\lambda$). In our framework, this scaling arises because the quantity that AdamW takes the EMA of (i.e. $q_{t}$ in Eq.6) scales with $1/\lambda$), but this also arises in Kosson et al. (2023). As we propose $\lambda\sim\text{fan\_in}$, this suggests that the learned weights (not the initial weights) scale as, $W_{ij}\sim 1/\lambda\sim 1/\text{fan\_in}$. Thankfully, this is precisely the scaling we need to ensure that,
$y_{i}$ $=\sum_{j=1}^{\text{fan\_in}}W_{ij}x_{j}$ (15)
does not blow up or shrink as the model size increases (specifically, $y_{i}\sim 1$). In particular, the learned weights, $W_{ij}$ are a sum of many updates, $\Delta W_{ij}$. Thus, following the same reasoning as that for Eq. (2), we have $y_{i}\sim W_{ij}\times\text{fan\_in}$. Thus, to ensure that $y_{i}\sim 1$, we need $W_{ij}\sim 1/\text{fan\_in}$, which is precisely what we get by following our proposed scaling.
5 Related work
Our EMA view of AdamW provides an explanation for several striking observations made in the literature.
First, the original AdamW paper (Loshchilov & Hutter, 2018) proposed a parameterization in terms of the learning rate, $\eta$ and $\gamma=\eta\lambda$ describing the weight decay. They made an empirical observation that $\eta$ and $\gamma$ were more decoupled than $\eta$ and $\lambda$ (in the sense that the optimal $\eta$ depended strongly on $\lambda$ and less strongly on $\gamma$). We are able to provide a theoretical understanding of their observation, by noting that $\gamma$ is related to our timescale, $\gamma=\nicefrac{{1}}{{\tau_{\text{iter}}}}$. We are therefore able to identify two separate processes one associated with $\eta$ and another associated with $\gamma=\nicefrac{{1}}{{\tau_{\text{iter}}}}$. First, $\eta$ describes the size of the initial updates relative to the random initialization, which is relevant close to initialization (see Appendix A for details). In contrast, $\gamma=\nicefrac{{1}}{{\tau_{\text{iter}}}}$ describes the EMA timescale, which is relevant as the model approaches convergence.
Second, this viewpoint suggests that the effects of the learning rate, $\eta$, on the magnitude of the updates relative to the random initialization should become less relevant than the EMA timescale as training proceeds (Appendix A). While this seems strange (of course the learning rate matters alot), there are some suggestions in the literature that indeed the EMA timescale may be more important in some settings. Specifically, Wortsman et al. (2024) ran proxies for large-scale LLM pretraining. They showed that if $\gamma=\eta\lambda=\nicefrac{{1}}{{\tau_{\text{iter}}}}$ is fixed, the final validation loss is relatively insensitive to the learning rate, $\eta$. In contrast, if $\lambda$ is fixed, the validation loss shows much more pronounced sensitivity to the learning rate, $\eta$. This makes sense if the key hyperparameter is $\tau_{\text{iter}}$, as modifying the learning rate, $\eta$, with $\lambda$ fixed implies $\tau_{\text{iter}}=1/(\eta\lambda)$ also changes.
Third, Lingle (2024) assessed the accuracy of µP in large-scale LLM pretraining. Remarkably, they found that the usual µP scaling of the optimal learning rate with model size broke down for AdamW. We provide an explanation and fix for this by noting that µP theory considers only the behaviour of the first few optimization steps, relative to the random initialization. The issue is that if the weight decay, $\lambda$, is fixed, then $\tau_{\text{iter}}$ changes with model size (as $\tau_{\text{iter}}$ depends on $\eta$, and $\eta$ changes with model size following the usual µP recommendations). We propose a fix for issue by proposing to strengthen weight decay as the model size increases, and validate the fix experimentally.
Finally, µP itself is related work (Yang et al., 2022), and is discussed extensively in the Background section. Additionally, the $\gamma=\eta\lambda$ parameterization is often used in SGD, where it is identified as the effective learning rate (Van Laarhoven, 2017; Zhang et al., 2019; Li & Arora, 2019; Li et al., 2020; Wan et al., 2021; Li et al., 2022), though this literature is not clear on whether those results transfer to Adam or AdamW. Finally, recent work (Busbridge et al., 2023) studies the scenarios where an EMA is used for averaging model weights, such as in self-supervised learning, and proposes hyperparameter scaling rules to achieve the same performance under different batch sizes. Our work differs in that we are interpreting AdamW as EMA over past updates rather than using explicitly using EMA for averaging weights.
6 Limitations
There are other update rules such as Lion (Chen et al., 2023) and Sophia (Liu et al., 2024) that use decoupled weight decay. We expect that our recommendations will transfer to those settings, though we have not investigated this explicitly as AdamW is by far and away the most popular method at present.
7 Conclusions
We showed that AdamW’s weight updates can be understood as an EMA. The critical EMA hyperparameter is the EMA timescale, $\tau_{\text{epoch}}$, and we can use $\tau_{\text{epoch}}$ (in combination with the learning rate, $\eta$) to set the weight decay. We gave guidelines for setting $\tau_{\text{epoch}}$, and argued that $\tau_{\text{epoch}}$ should remain fixed as we change model and dataset size. This implies that the optimal weight decay should decrease with dataset size and increase with model size, and we validated these predictions experimentally.
Appendix A Confirming the connection between AdamW and weight decay
The key insight from the EMA view on AdamW is that an EMA (Eq. 5), has only two choices, the EMA timescale, $\tau_{\text{iter}}$, and the initialization, $\text{ema}_{0}$. Here, we prove that there are similarly only two choices when using AdamW in a scale-invariant network (see Eq. 17). In particular, we show that for a carefully chosen initialization, the full trajectory of weights optimized by AdamW is controlled by $\tau_{\text{iter}}$ and the size of the first step, relative to the initial parameter scale $\rho=\eta/\sigma$.
Theorem 1.
Consider two AdamW optimizers with different learning rates and weight decays, and initialization scales, ($\eta,\lambda,\sigma$ vs. $\eta^{\prime},\lambda^{\prime},\sigma^{\prime}$). Take $w_{t}$ to be the parameters learned by the first optimizer after the $t$th optimization step, and $w_{t}^{\prime}$ to be the parameters learned by the second optimizer. These optimizers are initialized by
$w_{0}$ $=\sigma{\xi}$ (16a)
$w_{0}^{\prime}$ $=\sigma^{\prime}{\xi},$ (16b)
where $\boldsymbol{\xi}$ is random noise (e.g. IID Gaussian).
Consider a scale-invariant network, in the sense that multiplying the weights, $w$, by an arbitrary positive constant, $c>0$ gives the same output for all inputs, $x$,
$\operatorname{net}(x;w)$ $=\operatorname{net}(x;\tfrac{1}{c}w).$ (17)
We use the same EMA timescale,
$\nicefrac{{1}}{{\tau_{\text{iter}}}}=\eta\lambda=\eta^{\prime}\lambda^{\prime},$ (18)
and the same ratio of the first step, relative to the initial parameter scale,
$\rho=\tfrac{\eta}{\sigma}=\tfrac{\eta^{\prime}}{\sigma^{\prime}}.$ (19)
Now, the entire trajectory of weights for each optimizer is the same (up to a scalar multiplier,)
$w_{t}^{\prime}$ $=\tfrac{1}{c}w_{t},\forall t$ (20)
where
$c$ $=\tfrac{\sigma}{\sigma^{\prime}}=\tfrac{\eta}{\eta^{\prime}}=\tfrac{\lambda^{\prime}}{\lambda}.$ (21)
Thus, the network outputs are the same at all points along the trajectory,
$\operatorname{net}(x;w_{t})$ $=\operatorname{net}(x;w_{t}^{\prime}).$ (22)
A.1 Proof of Theorem 1
Here, we consider two AdamW training trajectories with hyperparameters $\eta,\lambda,\epsilon$ and $\eta^{\prime},\lambda^{\prime},\epsilon^{\prime}$. We prove that for a specific way of initializing the optimizer state, and if the optimizer hyperparameters are related by,
$\eta^{\prime}$ $=\tfrac{1}{c}\eta$ (23)
$\lambda^{\prime}$ $=c\lambda$ (24)
$\epsilon^{\prime}$ $=c\epsilon.$ (25)
then the full trajectories are the same. Notice that $\nicefrac{{1}}{{\tau_{\text{iter}}}}=\eta\lambda=\eta^{\prime}\lambda^{\prime}$ remains unchanged with this choice of hyperparameters.
We consider AdamW updates of the form,
$g_{t}$ $=\nabla\mathcal{L}(w_{t-1})$ (26a)
$m_{t}$ $=\beta_{1}m_{t-1}+(1-\beta_{1})g_{t}$ (26b)
$v_{t}$ $=\beta_{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}$ (26c)
$\hat{m}_{t}$ $=\frac{m_{t}}{1-\beta_{1}^{t}}$ (26d)
$\hat{v}_{t}$ $=\frac{v_{t}}{1-\beta_{2}^{t}}$ (26e)
$w_{t}$ $=(1-\lambda\eta)w_{t-1}+\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (26f)
(And the analogous equations for the primed optimizer). The optimizer state is the variables that persist across timesteps, namely $m$, $v$, and $w$. Specifically, the first optimizer has state $m_{t}$, $v_{t}$, and $w_{t}$ while the second optimizer has state $m^{\prime}_{t}$, $v^{\prime}_{t}$, and $w^{\prime}_{t}$. Our goal is to prove that for a scale-invariant network, as $\eta$ or $\lambda$ changes but $\nicefrac{{1}}{{\tau_{\text{iter}}}}=\eta\lambda$ is fixed, the states for the different optimizers are related by scale parameters,
$m^{\prime}_{t-1}$ $=cm_{t-1},$ (27a)
$v^{\prime}_{t-1}$ $=c^{2}v_{t-1},$ (27b)
$w^{\prime}_{t-1}$ $=\tfrac{1}{c}w_{t-1}.$ (27c)
where $c=\lambda^{\prime}/\lambda=\eta/\eta^{\prime}$. The argument proceeds inductively, by assuming that the scaling relations in Eq. (27) hold at timestep t1t-1, then proving that as a consequence they hold at timestep $t$.
We start by considering the updates for $m$ and $v$, which require the scaling of the gradients. In particular, assuming, that Eq. (27c) holds at timestep $t$, Appendix A.2 tells us that the relationship between gradients is,
$g^{\prime}_{t}=\nabla\mathcal{L}(w^{\prime}_{t-1})=c\nabla\mathcal{L}(w_{t-1})=cg_{t}.$ (28)
Thus, the updates for $m^{\prime}$ are,
$m^{\prime}_{t}$ $=\beta_{1}m^{\prime}_{t-1}+(1-\beta_{1})g^{\prime}_{t}.$ (29)
substituting Eq. (27a) and Eq. (28)
$m^{\prime}_{t}$ $=\beta_{1}cm_{t-1}+(1-\beta_{1})cg_{t}=c(\beta_{1}m_{t-1}+(1-\beta_{1})g_{t})=cm_{t}.$ (30)
So the relationship continues to hold for $m$ at timestep $t$. For $v$ updates,
$v^{\prime}_{t}$ $=\beta_{2}v^{\prime}_{t-1}+(1-\beta_{2})(g^{\prime})_{t}^{2}.$ (31)
substituting Eq. (27b) and Eq. (28),
$v^{\prime}_{t}$ $=\beta_{2}c^{2}v_{t-1}+(1-\beta_{2})cg_{t}^{2}=c^{2}\left(\beta_{2}c^{2}v_{t-1}+(1-\beta_{2})g_{t}^{2}\right)=c^{2}v_{t}$ (32)
So the relationship continues to hold for $v$ at timestep $t$.
Next, we consider $\hat{m}$ and $\hat{v}$, which are not in themselves state variables, as they can be computed directly from $m$ or $v$ at that timestep.
$\hat{m}^{\prime}_{t}$ $=\frac{1}{1-\beta_{1}^{t}}m^{\prime}_{t}=\frac{1}{1-\beta_{1}^{t}}cm_{t}=c\hat{m}_{t}$ (33a)
$\hat{v}^{\prime}_{t}$ $=\frac{1}{1-\beta_{1}^{t}}v^{\prime}_{t}=\frac{1}{1-\beta_{1}^{t}}c^{2}v_{t}=c^{2}\hat{v}_{t}.$ (33b)
So the scaling relationships for $\hat{m}$ and $\hat{v}$ are analogous to those for $m$ and $v$.
Finally, we consider the weight updates themselves. The weight updates for the two optimizers are,
$w_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w_{t-1}+\eta\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (34a)
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w^{\prime}_{t-1}+\eta^{\prime}\frac{\hat{m}^{\prime}_{t}}{\sqrt{\hat{v}^{\prime}_{t}}+\epsilon^{\prime}}$ (34b)
Substituting $\eta=\nicefrac{{1}}{{\tau_{\text{iter}}}}/\lambda$ and $\eta^{\prime}=\nicefrac{{1}}{{\tau_{\text{iter}}}}/\lambda^{\prime}$,
$w_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}+\epsilon}$ (35a)
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w^{\prime}_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{\lambda^{\prime}}\frac{\hat{m}^{\prime}_{t}}{\sqrt{\hat{v}^{\prime}_{t}}+\epsilon^{\prime}}$ (35b)
Now, substituting Eq. (25) and Eq. (24) into the form for $w_{t}^{\prime}$,
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w^{\prime}_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{c\lambda}\frac{c\hat{m}^{\prime}_{t}}{\sqrt{c^{2}\hat{v}^{\prime}_{t}}+c\epsilon}$ (36)
Substituting Eq. (27c) and Eq. (33),
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})\tfrac{1}{c}w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{c\lambda}\frac{c\hat{m}_{t}}{\sqrt{c^{2}\hat{v}_{t}}+c\epsilon}$ (37)
Cancelling $c$,
$w^{\prime}_{t}$ $=(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})\tfrac{1}{c}w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{c\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}}$ (38)
And pulling out $1/c$,
$w^{\prime}_{t}$ $=\tfrac{1}{c}\left((1-\nicefrac{{1}}{{\tau_{\text{iter}}}})w_{t-1}+\nicefrac{{1}}{{\tau_{\text{iter}}}}\frac{1}{\lambda}\frac{\hat{m}_{t}}{\sqrt{\hat{v}_{t}}}\right)=\tfrac{1}{c}w_{t}.$ (39)
As required.
Thus, we have shown that if the scaling relations in Eq. (27) hold at timestep t1t-1, they also hold at timestep $t$. It only remains to complete the inductive proof by showing that the scaling relations hold at initialization. The $m$ and $v$ state variables are usually initialized at zero, $m_{0}=v_{0}=0$, and the scaling relations of course hold at zero. However, networks are usually initialized with a fixed parameter scale which does not depend on $\lambda$. To ensure full scale-invariant training, we therefore need to rescale the parameter initializations, to ensure $w_{0}^{\prime}=\tfrac{1}{c}w_{0}=\tfrac{\eta^{\prime}}{\eta}w_{0}$.
A.2 Gradients increase as weights decrease in scale-invariant networks
Figure 5: An image of a 2D scale-invariant loss, clearly showing that the gradients get larger closer to the origin.
One important component of our proof is the notion that as weights decrease, the gradient of a scale-invariant loss increases (e.g. see Arora et al., 2019). The geometry is straightforward if you look at an image (Fig. 5). But to spell it out formally, consider a scale-invariant loss,
$\mathcal{L}(w)$ $=\mathcal{L}(u)$ (40)
where $w=au$. Now, take the partial derivative and use $\frac{\partial u_{i}}{\partial w_{i}}=1/a$,
$\frac{\partial\mathcal{L}(w)}{\partial w_{i}}$ $=\frac{\partial\mathcal{L}(u)}{\partial u_{i}}\frac{\partial u_{i}}{\partial w_{i}}=\frac{1}{a}\frac{\partial\mathcal{L}(u)}{\partial u_{i}}$ (41)
Thus, bigger $w$ (larger $a$) implies smaller gradients wrt $w$.
A.3 Empirical confirmation of Theorem 1
[…]
This suggests that standard network setups require considerable modification before training trajectories actually become invariant to $\eta$ for a fixed $\nicefrac{{1}}{{\tau_{\text{iter}}}}$ (or $\tau_{\text{epoch}}$). In particular, recall that Theorem 1 has two major assumptions about the model: 1. It needs to be scale-invariant (Eq. 17); 2. The initialization needs to be $\eta$-aware (Eq. 19). As such, we need to make three key modifications in order for the assumptions to hold:
  1. 1.
    The output weights are not scale-invariant in standard setups, as they are not usually followed by a normalization layer; we therefore introduced a normalization layer after these weights (Appendix A.3.1; row C in Fig. 67)
  2. 2.
    Normalization layers such as batchnorm have scale and bias parameters, which are usually not scale-invariant, as such we adopted a decoupled learning rate for these weights (Appendix A.3.3; row D in Fig. 67).
  3. 3.
    We introduced an initialization that depended on the learning rate, $\eta$ (Appendix A.3.2; row E in Fig. 67).
Combining all these modifications together, we found that networks with the same $\tau_{\text{epoch}}$ but different values of $\eta$ demonstrated exactly the same learning trajectories (Fig. 6F) and performance metrics (Fig. 7F).
We additionally considered ViT under a similar setting. Notice that ViT needs more extensive modifications to ensure full scale-invariance; full details are presented in Appendix. A.3.4. The performance vs. iteration and timescale for various versions of the model are presented in Fig. 8 and 9. Mirroring the ResNet results, we again found that, under the suggested modifications, the performance becomes invariant to $\eta$ under fixed $\tau_{\text{epoch}}$.
Figure 6: Proposed modifications decouple the training trajectory from $\eta$ under fixed $\tau_{\text{epoch}}$. We trained a ResNet under various learning rate $\eta$ (lines) and timescale $\tau_{\text{epoch}}$ (columns) and plotted the training loss against epochs. We considered standard vs. normalized output layer (row B vs. row C, Sec. A.3.1); using $\eta$ for batchnorm parameters vs. a separate decoupled learning rate (row C vs. row D, Sec. A.3.3); using standard initialization scale vs. $\eta$-dependent initialization (row C vs. row D, Sec. A.3.2; and lastly, combining all modifications (row F), which allows the trajectory to be independent of $\eta$ under a fixed timescale.
Figure 7: Under proposed modifications, the performance is only controlled by the timescale. We trained a ResNet-18 on CIFAR-10 and plot the metrics at the end of the optimization under different initial learning rates (lines) against the timescale (x-axis). Similar to Fig. 6, from row B to F, we considered different levels of modification to the network. From row B to E, while the optimal $\tau_{\text{epoch}}$ (marked by red crosses) lies in close range across different $\eta$, the exact performance still varies by $\eta$. Whereas in row F, after adopting all three modifications, the performance metrics become invariant to $\eta$ under the same $\tau_{\text{epoch}}$.
Figure 8: Proposed scale-invariant structure decouples the training trajectory from $\eta$ under fixed $\tau_{\text{epoch}}$ for ViT. We trained a ViT under various learning rates $\eta$ (lines) and timescale $\tau_{\text{epoch}}$ (columns) and plotted the training loss against epochs. We first considered standard model vs. model with the three modifications used in ResNet experiments (row B vs. row C, Sec. A.3.1A.3.3A.3.2), where the training loss traces still show discrepancy for different $\eta$. We then considered further modification (Sec. A.3.4, row D) to ensure scale-invariance (SI), which successfully decoupled the loss trajectory from $\eta$ under the same $\tau_{\text{epoch}}$.
[…]
See below for more details of the network modifications required to ensure that experimental results can mirror Theorem 1 in Sec. A.
A.3.1 Output-batchnorm for scale-invariant output weights
Most neural network weights/parameters are scale-invariant. This means that we can multiply the weights/ parameters by an arbitrary constant without changing the network output (Eq. 17). It turns out that most neural network weights/parameters are scale-invariant because they are followed by normalization such as batchnorm or layernorm, which will get rid of any change in scale. However, this is not true for the output layer, as the output layer is not usually followed by a normalization layer.
As such, we applied a “global batchnorm” layer after the output layer, i.e., upon the logits. In particular, denote the logits by $Z\in\mathbb{R}^{B\times C}$, where $B$ is the batchsize and $C$ is the number of classes, we first flatten (and reshape) $Z$ into a matrix of shape $BC\times 1$ and then feed the vector into a standard 1D batchnorm, which computes the mean and variance across all $BC$ elements. Finally, we reshape the output of batchnorm back to $B\times C$ and send it to the final softmax layer. The effect of output-batchnorm is shown in row C in Fig. 6 and 7.
A.3.2 Learning-rate-dependent initialization
Note that Theorem 1 requires the ratio between the initial learning rate and the initial parameter scale to be the same in order for two AdamW optimizers to show the same trajectory. To satisfy the condition in implementation, we fix $\rho=\eta/\sigma=10^{-3}$, so the initial variance depends on the learning rate, $\sigma=\eta/10^{-3}$.
Confirming this requirement, initializing the network weights in this way seemed to reduce the dependency on $\eta$ for larger values of $\eta$ under fixed $\tau_{\text{epoch}}$ (Fig. 7E).
However, it is worth emphasizing that, $\rho$ itself is another hyperparameter, and different $\rho$ will change the performance (indeed, if we fix the initial variance, then $\rho$ changes as we modify $\eta$).
A.3.3 Decoupling the learning rates for the batch/layernorm parameters
In most networks, there are some parameters that are fundamentally non-scale-invariant. These usually include the scale and bias parameters in normalization layers. All of the reasoning in Theorem 1 rely on network output being invariant to scale. So how do we optimize these non-scale-invariant parameters?
It turns out that we already do something different with these parameters. In particular, it is common practice [^3] We cannot find literature explicitly studying the effect of dropping weight decay for normalization layers’ parameters, however, its importance has been noticed in vision tasks, e.g. by Jia et al. (2018) and https://github.com/JiahuiYu/slimmable_networks/issues/15, as well as in language model training: https://github.com/karpathy/minGPT/blob/master/mingpt/model.py#L227 to optimize them using the same learning rate, $\eta$, as the other parameters, but to drop weight decay. This standard setting is depicted in Fig. 6E and Fig. 7E, and we can see that $\eta$ does still change the behavior of the optimizer.
The issue is that modifying $\eta$, changes the training trajectory of these non-scale-invariant parameters. The solution is therefore to “decouple” the learning rate for the non-scale-invariant parameters. In particular, we fix the initial learning rate for these non-scale-invariant parameters to $10^{-3}$ regardless of the learning rate $\eta$ for other scale-invariant parameters. Doing so gives Fig. 6F and Fig. 7F, which exhibit almost perfect decoupling, with performance almost entirely independent of $\eta$ for fixed $\tau_{\text{epoch}}$.
A.3.4 Scale-invariant ViT
In order to make ViT scale-invariant, we need more extensive modifications beyond simply adding output normalization layer. In particular, we added layernorm after all linear layers in the network, which includes
  • The embedding layer.
  • The query, key and value projection layers.
  • The output projection layer after the attention operation.
  • The linear layer following each attention block.
Appendix B The weights in an EMA
Using Eq. (5), and assuming $\operatorname{ema}_{0}=0$,
$\operatorname{ema}_{1}$ $=\nicefrac{{1}}{{\tau_{\text{iter}}}}q_{1}$ (42)
$\operatorname{ema}_{2}$ $=\nicefrac{{1}}{{\tau_{\text{iter}}}}\left((1-\nicefrac{{1}}{{\tau_{\text{iter}}}})q_{1}+q_{2}\right)$ (43)
$\operatorname{ema}_{3}$ $=\nicefrac{{1}}{{\tau_{\text{iter}}}}\left((1-\nicefrac{{1}}{{\tau_{\text{iter}}}})^{2}q_{1}+(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})q_{2}+q_{3}\right)$ (44)
Thus,
$\operatorname{ema}_{t}$ $=\nicefrac{{1}}{{\tau_{\text{iter}}}}\sum_{t^{\prime}=1}^{t}(1-\nicefrac{{1}}{{\tau_{\text{iter}}}})^{t-t^{\prime}}q_{t^{\prime}}.$ (45)
In deep learning settings, $\tau_{\text{iter}}$ is much larger than one (implying that we average over many iterations). In that setting, the first-order Taylor expansion is very accurate,
$1-\nicefrac{{1}}{{\tau_{\text{iter}}}}\approx e^{-\nicefrac{{1}}{{\tau_{\text{iter}}}}}.$ (46)
Thus,
$\operatorname{ema}_{t}$ $\approx\nicefrac{{1}}{{\tau_{\text{iter}}}}\sum_{t^{\prime}=1}^{t}(e^{-\nicefrac{{1}}{{\tau_{\text{iter}}}}})^{t-t^{\prime}}q_{t^{\prime}}=\nicefrac{{1}}{{\tau_{\text{iter}}}}\sum_{t^{\prime}=1}^{t}e^{-(t-t^{\prime})/\tau_{\text{iter}}}q_{t^{\prime}}.$ (47)
with exact equality as $\tau_{\text{iter}}$ approaches infinity.
Appendix C Additional experiment results
We additionally considered training a ViT on the 320K subset of ImageNet 32x32. In particular, we used a ViT with QK-layernorm similar to the one used in the main text, and we varied the width/size of the model by multiplying the number of the hidden dimensions and internal linear layers’ width with a factor $s$. Similar to the setting in the main text, we tested the width factor in $\{0.5,1.0,2.0\}$, used a fixed $\lambda_{\text{base}}=1.0$, and swept $\eta_{\text{base}}$ in $(2.5\times 10^{-6})\times 2^{i}$ with $i\in\{0,1,\dotsc,11\}$.
The results are presented in Fig. 10, where we plotted the performance after 50 epochs against $\eta_{\text{base}}$. We again considered the direct µP scaling (Eq. 11; top row) and our proposed scaling (Eq. 13; bottom row), and we can see that the standard scaling breaks the stability of optimal learning rate in terms of test metrics whereas our proposed scaling is much more stable.
Figure 10: AdamW breaks the learning rate scaling of µP on ViT. Similar to the setting in Fig. 4, here we trained a ViT with different width factors on the 320K subset of ImageNet 32x32 under the direct µP scaling (Eq. 11; top row) and our proposed scaling (Eq. 13; bottom row). The direct scaling breaks the transferability of optimal $\eta_{\text{base}}$ due to changing the timescale, whereas our scaling allows optimal stable $\eta_{\text{base}}$ across model sizes.
Appendix D Note on muP library
We used the mup library [^4] https://github.com/microsoft/mup from the authors of Yang et al. (2022) for varying-model-size experiments. In particular, for ResNet experiments in Fig. 3 and 4, we directly use the ResNet codebase under the examples folder together with the provided MuAdamW optimizer to run our experiments, as this library takes care of details such as the scaling of the learning rate and initialization for the input/output layers. For ViT experiments in Fig. 10, we manually constructed the required model shape file using the provided make_base_shapes function and used the provided MuReadout module as the classification head.
Importantly, this library does come with a decoupled keyword argument. Given the connections between $\tau_{\text{iter}}$ and the parameterization for AdamW originally proposed in the original “Decoupled Weight Decay Regularization” paper (Loshchilov & Hutter, 2018), you would have thought that you could implement our proposed scalings using decoupled=True. However, it turns out that as of writing, to get our proposed scaling for $\lambda$ (Eq. 13), you need to set decoupled=False [^5] https://github.com/microsoft/mup/blob/19814971934ef91dd546f88e913fc963e096d11c/mup/optim.py#L79. This may be fixed in the future, but in any case, if using the mup library, it is critical to check the mup source to see precisely what scalings you are getting for $\lambda$.
Appendix E Extended experiment setups
E.1 Model specification
For ResNet-18 experiments, we utilized the implementation from https://github.com/kuangliu/pytorch-cifar/. For both CIFAR-10 and ImageNet, we used random cropping and horizontal flip as data augmentation and we used cross-entropy as the loss function.
For ViT, we adopted the implementation from https://github.com/omihub777/ViT-CIFAR/tree/main. We also incorporated QK layernorm suggested by Dehghani et al. (2023) and Wortsman et al. (2024) in order to stabilize the training when swiping learning rates. The loss function is again chosen as cross-entropy loss. Additionally, when training on CIFAR-10, we follow the suggestions in the original codebase to use auto augmentation (Cubuk et al., 2019) and label smoothing (Müller et al., 2019) with $\alpha=0.1$ to alleviate overfitting. We indeed found these techniques crucial for reaching the level of test accuracy reported by the repo. For ImageNet experiments, we used label smoothing with standard data augmentation: random cropping and horizontal flip.
E.2 Hyperparameter range
In Sec. 4.1, for each given $\eta$, we considered, $\eta\lambda$ of
$\nicefrac{{1}}{{\tau_{\text{iter}}}}=\eta\lambda\in\{10^{-7},3\times 10^{-7},10^{-6},3\times 10^{-6},\ldots,3\times 10^{-2},10^{-1}\}.$ (48)
In Sec. 4.2, for the experiments in Fig. 2, we used $\lambda=(2.5\times 10^{-6})\times 2^{i}$ with $i\in\{0,1,\dotsc,15\}$.
In Sec. 4.3, for the experiments in Fig. 3, we used $\lambda_{\text{base}}=10^{-3}\times 2^{i}$ with $i\in\{0,1,\dotsc,11\}$.
For the CIFAR-10 experiments in Fig. 4A we used $\eta_{\text{base}}=(2.5\times 10^{-4})\times 4^{i}$ with $i\in\{0,1,\dotsc,5\}$.
For the ImageNet experiments in Fig. 4B we used $\eta_{\text{base}}=(2.5\times 10^{-6})\times 2^{i}$ with $i\in\{0,1,\dotsc,11\}$.
Appendix F Licenses